# Cell bulging and extrusion in a three-dimensional bubbly vertex model for curved epithelial sheets

### v1

@@ Abstract
Cell extrusion is an essential mechanism for controlling cell density in epithelial tissues. Another essential element of epithelia is curvature, which is required to achieve complex shapes, like in the lung or intestine. Here we introduce a three-dimensional bubbly vertex model to study the interplay between extrusion and curvature. We find a generic cellular bulging instability at topological defects which is much stronger than for standard vertex models. To clarify the underlying mechanism, we analytically calculate the energy landscape for a single bulging cell in a mean field tissue. We find that extrusion is energetically preferred at topological defects due to a redistribution of Gaussian curvature into the defect. We show that it can be driven both by a decrease in apico-basal tension or by contractile line tensions. We also show that extrusion can be suppressed by luminal pressure and interfacial bending rigidity, allowing for additional control of this important process. Our theory suggests that epithelial curvature naturally leads to cell bulging and extrusion because the interfacial curvature of individual cells at the defects strongly amplifies buckling effected by tissue-scale topological defects in elastic sheets.
@@ I Introduction
Epithelial tissues are close-packed and sealed cell layers that form the barriers between different compartments in the body, for example in the lung or the intestine. In order to regulate their cell density, they have to achieve a balance between cell division and cell death. Recently it has become clear that both processes are strongly connected to mechanics: cells divide more if they are stretched by neighboring cells at low cell density, and they stop dividing and die if the epithelium becomes too crowded [1, 2, 3]. In order to make sure that dying cells leave the tissue without disrupting its barrier function, they are expelled by extrusion, which means they are squeezed out of the epithelium without the generation of any voids [4, 5]. Extrusion is especially prominent in the intestine, where the epithelium renews itself every 3-5 days [6]. Here cells proliferate in the crypt regions and then move in a constant stream to the tip of villi, where they are extruded [7]. The exact mechanisms of extrusion are not clear, but one important element seems to be the presence of specialized actin structures on the lateral and basal sides [3]. For flat model epithelia, it has been shown that cell extrusion occurs preferentially at topological defects in the nematic orientation field of the cells [8, 9, 10].
@@
In general, cells in non-stratified epithelia tend to have hexagonal order, because this is the close-packed optimum in two dimensions, similar to the situation in sphere packings and foams. Hexagonal order is also dominant in curved epithelial sheets [11, 12, 13, 14]. However, now some pentagonal defects are required to generate curvature [12]. Due to Euler’s polyhedron theorem, 12 pentagonal defects are required to achieve the topology of a sphere, with the possibility of additional pentagonal-heptagonal pairs. Thus topological defects occur naturally in curved epithelia, while in flat cell layers, they arise rather a in stochastic manner from the movement of cells. Because epithelia are often curved due to their biological function, for example in the crypts and villi of the intestine, the question arises if they might use the naturally occurring topological defects to promote cell extrusion. While previous theoretical models have addressed such extrusion in a flat configuration focusing on fluid-like epithelia [15] or neglecting the three-dimensional nature of the problem [16], the coupling of curvature, defects, and shape instabilities has hardly been addressed before. Defects in such a setting have been linked to extrusion in flat epithelia on a substrate in phase field models of tissues [15] and to shape instabilities in vertex models [17], but the underlying mechanisms and their relation to curvature have not been addressed yet.
@@
Recently, we have reported on a morphological instability at topological disclination defects in curved hexagonal epithelia in the three-dimensional vertex model (VM) [18]. This buckling instability arises because spherical epithelia are effectively elastic shells, similar to viral capsids [19, 20] or buckyballs [21], for which such an instability is known since the pioneering work by Seung and Nelson [22]. However, the standard VM does not consider that cells might adapt their own interfacial curvature to their mechanical environment, which must be the first step toward bulging and extrusion. Here we investigate this aspect by using the bubbly vertex model (BVM). In contrast to the normal VM, which describes multicellular systems with flat interfaces, in the BVM the interfaces are curved [23, 24, 25]. Such interfacial curvature leverages the three-dimensional nature of the cellular system and allows for different mechanisms than the usual VM-approach to epithelia. Here we show that the BVM for curved epithelial sheets naturally explains why and how cells can extrude preferentially at topological defects.
@@
We start in Section II by showing with computer simulations that the curved interfaces in the BVM strongly amplify the buckling at defects in curved sheets if compared to the standard VM with flat interfaces. To understand the underlying instability theoretically, we consider the representative case of a shell with icosahedral symmetry in which interfacial curvature leads to a significant lowering of a previously described icosahedral buckling threshold [18, 19, 22]. In Section III we develop a continuum description of a defect cell in a mean field tissue, which allows us to analytically understand the underlying instability. We find that the possibility of curved interfaces results in a local decrease of the saddle splay modulus, thus leading to localization of Gaussian curvature in the defect. This analytical model captures the numerical shapes of icosahedral vertex model shells. Generalizing it in Section IV to shells without such strong symmetry properties, allows us to generally formulate a model for cell extrusion at pentagonal and hexagonal cells in a curved epithelial sheet. We find that both lowering the apico-basal surface tension and a supracellular actin ring may initiate bulging, while only the former allows for a transition of a partially extruded geometry toward full extrusion. Our model predicts a lower energy barrier for extrusion at topological (pentagonal) defects. Further, we show that this tendency can be reduced by increasing luminal pressure or rigidifying cellular interfaces. We close in Section V with a discussion of our results, including the relation to experimental observations.
@@ II.1 Definition of bubbly vertex model (BVM)
For the description of tissue mechanics and shape, we adopt the usual tension-based approach. We start from the regular three-dimensional vertex model for epithelial monolayers in which individual cells are described as bounded volumes with one apical, one basal and multiple lateral faces as boundaries. While in the classical vertex model these boundary faces are assumed to be flat by triangulating them with a passive interpolation point [26, 18], we explicitly want to consider the effect of possible interface curvature on the mechanics and thus introduce the bubbly vertex model, in which the interfaces are allowed to curve.
Figure 1: Cell bulging at topological defects in the bubbly vertex model. (a) Schematic depiction of the vertex model (VM) representation of a cell with apical, basal and lateral faces via an interpolation point of the polygonal faces, and of the bubbly vertex model (BVM) representation with curved interfaces. (b) Spherical VM shell after energy minimization with random initialization of cell centers. (c) The same shell in the BVM after minimization in Surface Evolver. Cell symmetry is color-coded. (d) Opening angles of the cells by symmetry for the VM and BVM, where the half angle is determined for each lateral interface (inset) and then averaged over lateral interfaces. Violin plots show distribution with the horizontal line marking the average value. Shells with $400$ cells and tensions $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}=0.9$, $\Gamma_{\mathrm{l}}=1$.
Each cell $i$ has one apical area $A_{a,i}$, one basal area $A_{b,i}$ and the integrated lateral area $A_{l,i}$, corresponding to the areas of the boundary faces. We assume the faces to have surface tensions $\Gamma_{(\cdot)}$, where we assume each cell to have the same surface tension for the apical, basal, and lateral faces. The total energy of the tissue is then described by the sum of the individual surface area contributions
@@
Due to the viscoelastic behavior of the cell cortex [27, 28], we neglect elastic in-plane contributions of the interfaces, as these will relax over sufficiently long time scales. For now, we also do not consider a bending energy contribution [25], making use of the analogy to soap films, which are purely driven by surface tensions. This can be justified by the small thickness of the cortex, which will lead to small bending rigidities in the case of large interfacial tensions. We will relax this assumption later.
@@
To differentiate between the model with and without membrane curvature, we call the former vertex model (VM) and the latter bubbly vertex model (BVM) throughout the rest of this article. Fig. 1(a) schematically depicts the two model classes for one hexagonal cell.
II.2 Instability of luminal interfaces at topological defects
@@
We consider a spherical epithelium with a randomized starting configuration. For simulations we use OrganoidChaste [29], a custom-written package for Chaste [30] implementing a three-dimensional vertex model for monolayers, which was used in Ref. [18] and is described in Appendix A. Cell centers are created using the random sequential adsorption mechanism [31] and then a Voronoi tesselation on the sphere to determine cell shape. Using a rate of active topological T1-transitions which is slowly reduced to zero, the energy is minimized to obtain a spherical starting configuration. The resulting minimal energy configuration is shown in Fig. 1(b).
@@
Next we minimized the energy in the BVM by using the VM configuration as starting configuration in the standard software Surface Evolver for energy minimization of surfaces under tension [32]. Again we minimize at constant cell volume and get the result shown in Fig. 1(c). As shown by the color code, for both the VM and the BVM, pentagons, hexagons and heptagons are distributed all over the surface. The spherical configuration necessitates curvature screening [33] through additional pairs of topological defects (pentagon-heptagon pairs) compared to an icosahedral configuration of a hexagonal lattice with 12 pentagons. While topological disclination defects indeed seem to be screened by the distribution of additional defect pairs in the VM, leading to an approximately spherical configuration, we observe an interesting instability in the BVM. Pentagonal cells tend to bulge outward, minimizing their luminal interfacial area, despite the fact that we do not observe a tissue-scale buckling instability. Fig. 1(d) demonstrates the dramatic difference between the two models: the opening angles of pentagonal cells are much larger in the BVM than in the VM, agreeing with the visual impression when comparing Figs. 1(b) and (c). In addition, we also see that the distribution of opening angles in the BVM is much broader with some cells even bending in the opposite direction. We conclude that the BVM allows for more adaption and variability in cell shape than the classical VM, with a particularly strong effect on the pentagonal defects.
II.3 Shift of icosahedral buckling threshold

### v2

# arXiv:2411.07141v2[physics.bio-ph] 27 Feb 2026

@@ Abstract
Cell extrusion is an essential mechanism for controlling cell density in epithelial tissues. Another essential element of epithelia is curvature, which is required to achieve complex shapes, like in the lung or intestine. Here we introduce a three-dimensional bubbly vertex model to study the interplay between extrusion and curvature. We find a generic cellular bulging instability at topological defects which is much stronger than for standard vertex models. Analyzing cell shapes in three-dimensional imaging data of spherical mouse colon organoids, we infer that pentagonal cells have an increased basal interfacial tension, suggesting that cells at topological defects react to the different force conditions. Using the bubbly vertex model, we show that such basal tensions stabilize against the predicted instability and result in better cell shape control than tissue-scale mechanisms such as lumen pressure and spontaneous curvature. Our theory suggests that epithelial curvature naturally leads to bulged and extrusion-like cell shapes because the interfacial curvature of individual cells at the defects strongly amplifies buckling effected by tissue-scale topological defects in elastic sheets. Our results highlight the complex interplay of forces across scales in three-dimensional tissue organization.
@@ I Introduction
Epithelial tissues are close-packed and sealed cell layers that form the barriers between different compartments in the body, for example in the lung or the intestine. In order to regulate their cell density, they have to achieve a balance between cell division and cell death. Recently it has become clear that both processes are strongly connected to mechanics: cells divide more if they are stretched by neighboring cells at low cell density, and they stop dividing and die if the epithelium becomes too crowded [1, 2, 3]. In order to make sure that cells leave the tissue without disrupting its barrier function, they are usually expelled by extrusion, which means they are squeezed out of the epithelium without the generation of any voids [4, 5]. Extrusion is especially prominent in the intestine, where the epithelium renews itself every 3-5 days [6]. Here cells proliferate in the crypt regions and then move in a constant stream to the tip of villi, where they are extruded [7]. Very recently, it has been shown that most cells extruded in the villus regions are still alive, which means that they mechanically do not differ much from their neighbors [8]. The exact mechanisms of extrusion are not clear, but one important element seems to be the presence of specialized actin structures on the lateral and basal sides [3, 8]. For flat model epithelia, it has been shown that cell extrusion occurs preferentially at topological defects in the nematic orientation field of the cells [9, 10, 11].
@@
In general, cells in non-stratified epithelia tend to have hexagonal order, because this is the close-packed optimum in two dimensions, similar to the situation in sphere packings and foams. Hexagonal order is also dominant in curved epithelial sheets [12, 13, 14, 15]. However, now some pentagonal defects are required to generate curvature [13]. Due to Euler’s polyhedron theorem, 12 pentagonal defects are required to achieve the topology of a sphere, with the possibility of additional pentagonal-heptagonal pairs. Thus topological defects occur naturally in curved epithelia, while in flat cell layers, they arise rather in a stochastic manner from the movement of cells. Because epithelia are often curved due to their biological function, for example in the crypts and villi of the intestine, the question arises if they might use the naturally occurring topological defects to promote cell extrusion. While previous theoretical models have addressed such extrusion in a flat configuration focusing on fluid-like epithelia [16] or neglecting the three-dimensional nature of the problem [17], the coupling of curvature, defects, and shape instabilities has hardly been addressed before. Defects in such a setting have been linked to extrusion in flat epithelia on a substrate in phase field models of tissues [16] and to shape instabilities in vertex models [18], but the underlying mechanisms and their relation to curvature have not been addressed yet.
@@
Recently, we have reported on a morphological instability at topological disclination defects in curved hexagonal epithelia in the three-dimensional vertex model (VM) [19]. This buckling instability arises because spherical epithelia are effectively elastic shells, similar to viral capsids [20, 21] or buckyballs [22], for which such an instability is known since the pioneering work by Seung and Nelson [23]. However, the standard VM does not consider that cells might adapt their own interfacial curvature to their mechanical environment. Here we investigate this aspect by using the theoretical bubbly vertex model (BVM) and by analyzing experimentally observed cell shapes in spherical mouse colon organoids. In contrast to the normal VM, which describes multicellular systems with flat interfaces, in the BVM the interfaces are curved [24, 25, 26, 27, 28]. Such interfacial curvature leverages the three-dimensional nature of the cellular system and allows for different mechanisms than the usual VM-approach to epithelia. On the experimental side, colon organoids are particularly suited to address the coupling of mechanics, cell shape and topology, as they form spherical epithelial monolayers. They are derived from stem cells gathered from colonic crypts, recapitulate developmental processes of the colon and thus serve as important model system for development and disease [29].
@@
We start in Section II by showing with computer simulations that the curved interfaces in the BVM strongly amplify the buckling at defects in curved sheets if compared to the standard VM with flat interfaces. Through segmentation of 3D microscopy data of mouse colon organoids with the subsequent inference of interfacial tensions, we find that cellular shapes are more regular in spherical organoids than theoretically expected. We also infer that outer (basal) tension are increased at pentagonal cells, suggesting that cells react to the changed force conditions at topological defects. In the BVM, we find such increased basal tensions to stabilize against the predicted instability and to yield full control over cellular bulging. To understand the underlying instability theoretically, we consider the representative case of a shell with icosahedral symmetry in which interfacial curvature leads to a significant lowering of a previously described icosahedral buckling threshold [19, 20, 23]. In Section III we develop a continuum description of a defect cell in a mean field tissue, allowing us to extend our model to describe and study extrusion-like cell geometries. We find that the possibility of curved interfaces results in a local decrease of the saddle splay modulus, thus leading to localization of Gaussian curvature in the defect. We find that lowering the apico-basal surface tension and a supracellular actin ring may initiate bulging, while only the former allows for a transition of a partially extruded geometry toward full extrusion. Our model predicts a lower energy barrier for extrusion at topological (pentagonal) defects. We close in Section IV with a discussion of our results.
@@ II.1 Definition of bubbly vertex model (BVM)
For the description of tissue mechanics and shape, we adopt the usual tension-based approach. We start from the regular three-dimensional vertex model for epithelial monolayers in which individual cells are described as bounded volumes with one apical, one basal and multiple lateral faces as boundaries. While in the classical vertex model these boundary faces are assumed to be flat by triangulating them with a passive interpolation point [30, 19], we explicitly want to consider the effect of possible interface curvature on the mechanics and thus introduce the bubbly vertex model, in which the interfaces are allowed to curve [24, 25, 26, 27, 28], cf. Fig. 1(a).
Figure 1: Cell bulging at topological defects in the bubbly vertex model. (a) Schematic depiction of the vertex model (VM) representation of a cell with apical, basal and lateral faces via an interpolation point of the polygonal faces, and of the bubbly vertex model (BVM) representation with curved interfaces. (b) Spherical VM and BVM shell after energy minimization with random initialization of cell centers. BVM shell is after minimization in Surface Evolver. Number of cell neighbors is color-coded. (c) Opening angles of the cells by neighbor number for the VM and BVM, where the half angle is determined for each lateral interface (inset) and then averaged over lateral interfaces. Violin plots show distribution with the horizontal line marking the mean value. Tensions $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}=0.9$, $\Gamma_{\mathrm{l}}=1$. (d) Mean opening angles of cells with different neighbor numbers (color-coded; red pentagons, black hexagons, blue heptagons) in different BVM-shells with random initial cell configurations ($N=6$) and varying apical and basal tensions $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}$ and $\Gamma_{\mathrm{l}}=1$. All shells with $400$ cells.
Each cell $i$ has one apical area $A_{a,i}$, one basal area $A_{b,i}$ and the integrated lateral area $A_{l,i}$, corresponding to the areas of the boundary faces. We assume the faces to have surface tensions $\Gamma_{\mathrm{a}}$, $\Gamma_{\mathrm{b}}$ and $\Gamma_{\mathrm{l}}$ for the apical, basal, and lateral faces, respectively, and that their values are the same for all cells (no dependance on $i$). The total energy of the tissue is then described by the sum of the individual surface area contributions
@@
Due to the fluid viscoelastic behavior of the cell cortex [31, 32], we neglect elastic in-plane contributions of the interfaces, as these will relax over sufficiently long time scales. For now, we also do not consider a bending energy contribution [26], making use of the analogy to soap films, which are purely driven by surface tensions. This can be justified by the small thickness of the cortex, which will lead to small bending rigidities in the case of large interfacial tensions. We will relax this assumption later.
@@
To differentiate between the model with and without membrane curvature, we call the former vertex model (VM) and the latter bubbly vertex model (BVM) throughout the rest of this article.
II.2 Instability at topological defects in the BVM
@@
We consider a spherical epithelium with a randomized starting configuration. For simulations we use OrganoidChaste [33], a custom-written package for Chaste [34] implementing a three-dimensional vertex model for monolayers, which was used in Ref. [19] and is described in Appendix A. Cell centers are created using the random sequential adsorption mechanism [35] and then a Voronoi tesselation on the sphere to determine cell shape. Using a rate of active topological T1-transitions which is slowly reduced to zero, the energy is minimized to obtain a spherical starting configuration. This process is performed in the VM to obtain shapes with low energy and a representative minimal energy configuration is shown in Fig. 1(b). This topology is then fixed for the comparison between VM and BVM shapes, as here we do not consider tissue dynamics on a larger time scale.
@@
Next we minimized the energy in the BVM by using the VM configuration as starting configuration in the standard software Surface Evolver for energy minimization of surfaces under tension [36]. Again we minimize at constant cell volume and get the result shown in Fig. 1(b). As shown by the color code, for both the VM and the BVM, pentagons, hexagons and heptagons are distributed all over the surface. The spherical configuration necessitates curvature screening [37] through additional pairs of topological defects (pentagon-heptagon pairs) compared to an icosahedral configuration of a hexagonal lattice with 12 pentagons. While topological disclination defects indeed seem to be screened by the distribution of additional defect pairs in the VM, leading to an approximately spherical configuration, we observe an interesting instability in the BVM. Pentagonal cells tend to bulge outward, minimizing their luminal interfacial area, despite the fact that we do not observe a tissue-scale buckling instability. Fig. 1(c) demonstrates the dramatic difference between the two models: the opening angles of pentagonal cells are much larger in the BVM than in the VM, agreeing with the visual impression when comparing shells in Fig. 1(b). In addition, we also see that the distribution of opening angles in the BVM is much broader with some cells even bending in the opposite direction. We conclude that the BVM allows for more adaptation and variability in cell shape than the classical VM, with a particularly strong effect on the pentagonal defects. We verified the generality of this result, by sampling many random configurations for shells with the same number of cells and for different apico-basal tensions $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}$, effectively considering different shell sizes with respect to the monolayer thickness, cf. Fig. 1(d).
II.3 Cell shapes at topological defects in mouse colon organoids
Figure 2: Experimentally observed cell shapes at topological defects in mouse colon (mCol) organoids. (a) Microscopy image of spherical mouse colon organoid. Green is membrane ($\beta$-catenin) and magenta is F-actin. The luminal side is apical (high actin) and the outer side basal (facing the matrix). (b) The organoid is segmented in 3D and the number of neighbors (color-coded) is determined from the basal polygonal edges. (c) For this organoid the opening angles $\delta/2$, averaged over all lateral interfaces, vary for different neighbor numbers. Each dot represents one cell, the bars depict the means. (d) The relative angle increase at pentagonal and heptagonal cells, compared to hexagonal cells. Every data point is one organoid ($N=22$). (e) Microscopy image of organoid treated with Latrunculin A (LatA). (f) Examples of pentagonal and hexagonal cells as segmented in 3D. (g) Opening angles for organoid treated with LatA. (h) Comparing the relative angle increase $\rho$ from wildtype ($N=13$) and LatA-treated samples ($N=17$), we find larger effects at topological defects with LatA. Scale bars in (a) and (e) are $20\,\mu\mathrm{m}$. Stars indicate: p<0.05*\ p<0.05, p<0.01**\ p<0.01, p<0.001***\ p<0.001. Tests against 0 in (d) with Wilcoxon and against each other in (h) with Mann-Whitney-U tests.
Mouse colon (mCol) organoids are spherical epithelia derived from single stem cells [29, 38]. Fig. 2(a) shows a typical microscopy image. We have grown these organoids in extracellular matrix (basement membrane extract) from mechanically dissociated (split) closed organoids and fixed, stained and imaged them at day 5 post-splitting. The luminal (inner) side of the organoids corresponds to the apical side of the cells, as demonstrated by the actin staining in Fig. 2(a). The outer side corresponds to the basal side of the cells and has been growing against the matrix. The three-dimensional images were segmented, as shown in Fig. 2(b), using a previously described pipeline that we have developed in the context of force inference [39] (see Appendices B and C for details). Similarly to the BVM, we determined the (half) cellular opening angles $\delta/2$ of the mCol cells as a function of their number of neighbors. Contrary to what we expect from the BVM, we find that in some organoids the opening angles of pentagonal cells are smaller than for hexagons and vice versa for heptagons (Fig. 2(c) for one organoid). To allow for a comparison across organoids, we determined the relative increase of cell-averaged pentagonal ($n=5$) and heptagonal ($n=7$) opening angles, $\langle\delta/2\rangle_{n}$, compared to the angles averaged over hexagonal cells, $\langle\delta/2\rangle_{6}$, i.e.
$\rho(n)=\frac{\langle\delta/2\rangle_{n}-\langle\delta/2\rangle_{6}}{\langle\delta/2\rangle_{6}}.$ (2)
The result is shown in Fig. 2(d). While we find this angle increase in pentagons to be larger than zero in some organoids, consistent with the bulging instability, it is significantly lower than zero on the level of many organoids. In heptagonal cells, this difference is not significant.
The experimental data show that cell shape is indeed significantly different at topological defects, albeit oppositely to the bulging instability predicted by the BVM. As the actomyosin systems has a stabilizing function through both the formation of an apical (luminal) actin belt and interfacial contractility, we treated the mCol organoids with Latrunculin A, an inhibitor of actin polymerization, effectively leading to actin depolymerization. While overall organoid morphology was not disrupted (Fig. 2(e)), we found a signature of the bulging instability in multiple treated organoids (Fig. 2(f,g)). Indeed, looking at the relative angle increase $\rho(n)$, we found increased angles at pentagonal cells in treated organoid populations, hinting at a potentially stabilizing role of the actomyosin system against the bulging instability at disclination defects (Fig. 2(h)). In heptagonal cells we do not see a significant difference between treated and untreated organoids, which could be explained by the smaller relevance of this instability here due to the lack of saddle-like structures in these spherical organoids, cf. Fig. 1(c).
Figure 3: Surface tension inference in mouse colon (mCol) organoids. (a) Assuming force balance, the relative tensions can be inferred from the dihedral angles of the interfaces. For the reconstructed organoid the outer (basal) tensions are shown. (b,c) For each wildtype (WT) (b) and Latrunculin A treated (LatA) (c) organoid we determine the difference of the averaged inferred tensions $\hat{\Gamma}$ for $n$-gons from hexagons for apical (a) and basal (b) interfaces. The averaged cell-specific apico-basal tension difference is also shown (a-b). (d) Pentagonal basal tension deviations versus relative angle increase of pentagons compared to hexagons $\rho(5)$ for WT and LatA samples. Solid line is linear regression, which intersects a tension deviation of $0.0$ at $\rho(5)\approx 0.17$. (e) Relative mid-plane area deviations of pentagons (red) and heptagons (blue) from hexagons for WT and LatA organoids. (f) Pentagonal basal tension deviations versus relative area deviations of pentagons for WT and LatA samples. Solid line is linear regression, which intersects a tension deviation of 0.0 at $\left(\langle A\rangle_{5}-\langle A\rangle_{6}\right)/\langle{A}\rangle_{6}=-0.015$. Pearson correlation coefficient in (d,f) is $\rho_{\mathrm{c}}$. Stars indicate: p<0.05*\ p<0.05, p<0.01**\ p<0.01, p<0.001***\ p<0.001. Tests against 0 in (b,c,e) with Wilcoxon test and against each other in (e) with Mann-Whitney-U-test.
Cellullar shapes display more regular opening angles than predicted by the BVM, suggesting more controlled force generation. In particular, mechanosensing and an active compensation mechanism may stabilize shape. In order to test this, we determined the interfacial tensions of the individual apical, basal and lateral interfaces, denoted with $\hat{\Gamma}$, from the microscopy images. Assuming mechanical equilibrium, interfacial tensions lead to forces normal to tri-interfacial junctions [39, 40, 41, 42, 43]. Using our recently developed force inference method [39], we inferred the underlying relative tensions from the interfacial dihedral angles in the segmented images (cf. Fig. 3(a), see Appendix B for details). Figures 3(b,c) display the different average tensions of different interface types at pentagonal and heptagonal cells, $\langle\hat{\Gamma}\rangle_{n}$, from the respective averages in hexagonal cells, $\langle\hat{\Gamma}\rangle_{6}$ for wildtype (WT) and Latrunculin A (LatA) treated cells. Apical tensions are not different in pentagonal and heptagonal cells, but basal tensions are higher (lower) at pentagonal (heptagonal) cells for WT cells. Treatment of LatA lowers this basal tension, as could be expected from the increased opening angles. The basal tension increase at pentagonal cells is negatively correlated with the relative angle increase at pentagonal cells (correlation $\rho_{\mathrm{c}}=-0.29$ with p-value $p=0.07$), supporting the notion of a complex balance between opening angles and basal tension. Linear regression on both the WT and LatA data suggests a relative angle increase of approximately $17\%$ if the pentagonal basal tension is not different from the hexagonal one (Fig. 3(d)). This indeed hints at an active compensation effect being responsible for the shape difference.
To explore potential mechanisms by which cells could adapt dynamically to their topological neighborhood, we determined the mid-plane areas as the ratio of cellular volume and height, $V_{\mathrm{cell}}/h$ (cf. Fig. 3(e)). Similarly to Lewi’s law in two-dimensional epithelia and foams which states that area increases with the number of neighbors [44, 45, 46], we find approximately $10\%$ smaller pentagonal and $15\%$ larger heptagonal areas. This enables a direct feedback of tissue topology to the cell and could be linked to mechanosensing, similarly as in density sensing in intestinal epithelia [1, 47]. Looking at the relationship of pentagonal basal tensions and areas, we find a negative correlation $\rho_{\mathrm{c}}=-0.34$ (p-value $p=0.03$): for smaller average pentagonal areas we find increased basal tensions (Fig. 3(f)). Linear regression suggests that this tension increase vanishes as areas become comparable to hexagonal areas. This is consistent with an area-driven feedback via mechanosensing facilitating a topology-dependent stabilization. Nonetheless, the underlying bulging instability is not as prominent in mouse colon organoids, implying the need for stabilization mechanisms also in our theoretical description.
II.4 Predictions of BVM for different shape stabilization mechanisms
Figure 4: Modulation of instability via different mechanisms explored through the BVM. (a) The shell from Fig. 1 simulated with additional effects: pressurized with $V/V_{\mathrm{VM}}=2.0$, rigidified with $\kappa_{\mathrm{if}}=0.7$, and pentagonal basal (outer) defect tension of $\Gamma_{\mathrm{b}}^{(5)}=0.957$. (b) Opening angles of pentagonal (red), hexagonal (black) and heptagonal cells (blue) as a function of the number of simulated cells. (c) Opening angles for different apico-basal tensions differences $\Delta\Gamma=\Gamma_{\mathrm{a}}-\Gamma_{\mathrm{b}}$, corresponding to different spontaneous curvatures. (d) Opening angles for different luminal volume constraints, where $V/V_{\mathrm{VM}}\approx 1$ corresponds to the case without a luminal volume constraint. (e) Opening angles for rigidified shells different interfacial bending rigidities $\kappa_{\mathrm{if}}$, where $\kappa_{\mathrm{if}}=0$ is the regular BVM. (f) Dependence on the basal apical tensions $\Gamma^{(5)}=\Gamma_{\mathrm{a}}^{(5)}=\Gamma_{\mathrm{b}}^{(5)}$ of pentagonal cells, with unchanged tensions in other cells. (g) Dependence on basal (outer) tension $\Gamma_{\mathrm{b}}^{(5)}$ of pentagonal cells which is changed compared to the outer tension of the other cells. All other tensions are unchanged compared to the reference. Markers are average angles and shaded regions display $5\%$ to $95\%$ quantiles. Single red lines in (f) and (g) are individual pentagonal cells and dotted vertical lines mark the reference case.
Our experiments have revealed a coupling of tissue topology to cell shape and force generation. In particular, the data hint at an active basal stabilization mechanism that controls cellular shapes at topological defects (pentagons) in organoids. This agrees with a very recent experimental report that basal contractility prevents extrusion in crypt regions [8], but it is challenging to identify the exact mechanisms given the complex situation in organoids. In order to better understand the physical processes that might be involved in cell extrusion and shape stabilization, we turned back to the BVM. Fig. 4 depicts the role of different tissue and cell properties on the opening angles at defects in simulated BVM-shells. Fig 4(a) displays exemplary shell shapes in modified BVM-shells. For increasing shell size, we find that the effect of the topological defects on the opening angles decreases (Fig. 4(b)). In particular, for large shell sizes we see that some pentagonal cells also display negative opening angles, which is effected by cells bulging into the lumen as a consequence of the reduced background curvature. Introducing spontaneous tissue curvature via an apico-basal tension difference $\Delta\Gamma=\Gamma_{\mathrm{a}}-\Gamma_{\mathrm{b}}$, modulates the opening angles slightly, but cannot fully stabilize cell shape (Fig. 4(c)), even for a spontaneous curvature radius of hexagonal cells which corresponds to the shell radius, here $\Delta\Gamma\approx 0.14$. In the following we will therefore usually neglect spontaneous curvature, i.e., consider $\Gamma\equiv\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}$. Figure 4(d) shows a shell whose luminal volume, $V$, was prescribed to be a multiple of the luminal volume of the relaxed VM shell, $V_{\mathrm{VM}}$, corresponding to an internal luminal pressure. Internal pressure is able to unfold pentagonal defect cells, cf. Fig. 4(a). Contrary to the different opening angles for different number of neighbors, cf. Fig. 1(c), increased pressure leads to convergence of the average opening angles of pentagons, hexagons and heptagons, see Fig. 4(d). In addition, we also see that the distribution of opening angles in pressurized BVM-shells is less broad, more closely resembling the classical VM-case. This means that pressure can be used to suppress the buckling instability caused by the surface tensions. To estimate the experimental level of pressure-mediated stabilization, we calculated the theoretically expected aspect ratio [39, 19] of hexagonal cells in the limit of flat interfaces (the classical VM) from inferred tensions (cf. Fig. 3(a); median relative tensions $\Gamma_{\mathrm{a}}\approx 1.02$, $\Gamma_{\mathrm{b}}\approx 1.06$, $\Gamma_{\mathrm{l}}\approx 0.97$) to be $\psi_{\mathrm{theo}}=h/A^{1/2}=2^{1/2}3^{-1/4}(\Gamma_{\mathrm{a}}+\Gamma_{\mathrm{b}})/\Gamma_{\mathrm{l}}\approx 2.3$ with cell height $h$. Experimentally we find $\psi_{\mathrm{exp}}=h^{3/2}V_{\mathrm{cell}}^{-1/2}\approx 1.3$ yielding an experimental to theoretical shape ratio of $\psi_{\mathrm{exp}}/\psi_{\mathrm{theo}}\approx 0.6$, which is consistent with a luminal pressure stretching the epithelium. Comparing this ratio to the pressurized theoretical BVM case, we find similar values for a luminal volume of $V/V_{\mathrm{VM}}\approx 2.25$ ($\psi_{\mathrm{exp}}/\psi_{\mathrm{theo}}=0.62)$, where cells have not fully unfolded. If we take average surface tensions of intestinal organoids from the literature, which are of the order $\langle\Gamma\rangle\approx 5\,\mathrm{nN}\,\mu\mathrm{m}^{-1}$ [48], we can roughly estimate the luminal pressure in the limit of flat interfaces (the classical VM) to be $P\approx 440\,\mathrm{Pa}$ (see Appendix D for details). In our case, these pressures alone cannot fully explain smaller opening angles of pentagonal cells as observed experimentally.
To highlight the role of interface curvature, we added an interfacial bending energy to the BVM-minimization for each interface $\Omega_{\mathrm{if}}$
with interfacial bending rigidity $\kappa_{\mathrm{if}}$, and (total) mean curvature of the interface $H_{\mathrm{if}}$. This model penalizes interfacial curvature energetically. Fig. 4(e) shows the opening angles, whose values quickly approach each other as $\kappa_{\mathrm{if}}$ is increased. However, rigidifying the interfaces does not achieve as much convergence of the opening angles toward each other as luminal pressure. In simulations we find the cells to have less curved interfaces, interpolating between the VM and the BVM, but this outcome does not agree with experimental observations, cf. Fig. 2.
Motivated by our experiments, we further changed the apical (inner) and basal (outer) tensions of the pentagonal defect cells to be different to the other cells to introduce topology-dependent variability (Fig. 4(f)). In this case, we see that increasing the defect tension leads to a reduction of pentagonal opening angles, where a transition through a shape consistent with hexagons toward negative opening angles occurs. Note that reducing the tension can drastically increase the pentagonal opening angle, already hinting at a potential extrusion mechanism through a tension decrease.
Fig. 4(g) shows the effect of increased outer (basal) tensions at pentagonal cells, as found experimentally in the mCol-organoids. We also assumed a non-vanishing luminal pressure, as this is likely to reflect the experimental situation. We find that the cells transition through a state consistent with hexagonal cells toward a state with approximately opposite opening angle, implying full cell shape control, cf. shell in Fig. 4(a). Note that the transition point is cell-specific, as it likely depends on the details of the neighborhood topology. This suggests a potential feedback mechanisms between local topology (and induced stresses) and interfacial tensions.
Beyond surface tensions, we also considered an increase in global basal line tensions, and found that it too can stabilize cell shapes, see Appendix E. The basal stabilization mechanism observed in experiments thus proves to enable reliable control over cell shape.
II.5 Shift of icosahedral buckling threshold